a 0.7650g of sample clay containing 20.50 moisture gave a precipitate potassium perchlorate weighing 0.3822g what is the percentage of k2o in clay on a dry basis
1
Expert's answer
2020-05-06T13:58:29-0400

The dry weight in the clay sample equals:

mdry = mtotal - mmoisture

where mdry - dry weight, mtotal - total weight, mmoisture - wet weight

As mmoisture = mtotal × 20.50 % = 0.7650 g × 20.50% = 0.157 g

From here:

mdry = mtotal - mmoisture = 0.7650 g - 0.157 g = 0.608 g

%w = m / mdry × 100%

where %w - weight percentage, m - K2O weight, mdry - dry weight

From here:

%w = (0.3822 / 0.608 g) × 100% = 63%


Answer: 63%